18 drecreased by 3 times d
Found 2 solutions by haileytucki, Jk22:
Answer by haileytucki(390)

18-3*d=
15*d
15d

Answer by Jk22(389)

18-3d (operation priorities : times before sub)
Wrong : (18-3)*d=15*d (order given by text reading order)